OpenStudy (anonymous):

Inequalities question: Sketch the region of the in the x-y plane within which the inequalities y>(x-2)^2 and y

OpenStudy (campbell_st):

just use this site, it makes life easy to graph things https://www.desmos.com/calculator

OpenStudy (campbell_st):

so then find the points of intersection a and b then its \[\int\limits_{a}^{b} (x) dx - \int\limits_{a}^{b} (x^2) dx\] hope it helps

OpenStudy (anonymous):

i have done the graph but i don't know the limits of integration

OpenStudy (anonymous):

as its two inequalities how should i find the boundaries?

OpenStudy (anonymous):

or what is a and b in the equation?

OpenStudy (campbell_st):

so what values of x do the 2 curves intersect at..?

OpenStudy (campbell_st):

you should be able to tell that from the graph

OpenStudy (anonymous):

i have solved it as an equation y=x and y=(x-2)^2 and the answer i get is -1 and 4 which i put as the boundaries

OpenStudy (campbell_st):

wow... I would have thought the curves intersect at x = 1 and x = 4

OpenStudy (anonymous):